OpenStudy (anonymous):

Jamal and his friends are conducting a car wash to raise money for a school play. Over the first weekend they raised $336.00, which was 80% of what they set out to earn. If they collect $21 for each car they wash, how many more cars must they wash? Can someone help me?

OpenStudy (hyperlunatic):

helping

OpenStudy (anonymous):

So first thing, you would divide 336 by 0.80 which gives you 420

OpenStudy (anonymous):

Thats how much they set out to earn.

OpenStudy (anonymous):

The equation would be 21x + 336 = 420

OpenStudy (anonymous):

Solve for x and done

OpenStudy (hyperlunatic):

now thats where you are wrong

OpenStudy (anonymous):

Explain please

OpenStudy (hyperlunatic):

I am now

OpenStudy (anonymous):

21x = 420 - 336

OpenStudy (anonymous):

21x = 84

OpenStudy (anonymous):

x = 84/21 or 4 more cars

OpenStudy (hyperlunatic):

ok, what you are figuring out is right, but that is not what the question asks

OpenStudy (anonymous):

x is the number of cars here

OpenStudy (hyperlunatic):

you need to find: x= 420/21

OpenStudy (anonymous):

@hyperlunatic that is what the question is asking for, not too sure what you're trying to get at but it will most likely just confuse @kayleedilbeck

OpenStudy (anonymous):

no because it says HOW MANY MORE CARS MUST THEY WASH

OpenStudy (anonymous):

They've already collected 336 of the 420

OpenStudy (anonymous):

they earn 21 for every car

OpenStudy (anonymous):

21x= 420 - 336

OpenStudy (anonymous):

Yeah the 420 answer confused me. I put in the number 4 for how many more cars are needed to be washed and that was the answer but not really sure how you got 4

OpenStudy (anonymous):

21x = 84

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

no that was just step 1 @kayleedilbeck

OpenStudy (anonymous):

Sorry for the confusion

OpenStudy (anonymous):

It's okay, thank you!